The length of a rectangle is 7 more than the width and the area is 98 sqft. what is the length and width of the rectangle?
Answer by TimothyLamb(4379)

L = w + 7
Lw = 98
---
Lw = 98
(w + 7)w = 98
ww + 7w = 98
ww + 7w - 98 = 0
---
the above quadratic equation is in standard form, with a=1, b=7, and c=-98
---
to solve the quadratic equation, plug this:
1 7 -98
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
7
-14
---
the negative root doesn't make sense as a length, so use the positive root:
---
w = 7 feet
L = 14 feet
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php